# Simulating Work Extraction in a Dinuclear Quantum Battery Using a Variational Quantum Algorithm

### v1

@@ III.1 Variational quantum thermalizer and protocol of optimal work extraction
In this work, we employ the Variational Quantum Thermalizer (VQT) algorithm Verdon et al. (2019); das Neves Silva et al. (2024) to approximate the thermal states of the Hamiltonian described in Eq. (1) das Neves Silva et al. (2024). To approximate the thermal states, we utilize a probabilistic quantum ansatz Verdon et al. (2019) $\rho_{\theta,\phi}$, parameterized by variables $\theta$ and $\phi$. These parameters define the structure of the quantum circuit and the corresponding probabilistic distribution, allowing the ansatz to represent mixed quantum states. Specifically, $\phi$ parameterizes the unitary transformations applied to the quantum circuit. At the same time, $\theta$ governs the distribution of probabilities responsible for preparing the initial states $\rho^{(i)}_{\theta}$, before being processed by the circuit parameterized by $\phi$.
The target thermal state is approximated by minimizing the cost function $\mathcal{L}_{\theta,\phi}$ given by the quantum relative free energy Verdon et al. (2019) of the ansatz for the target Hamiltonian:
$\mathcal{L}_{\theta,\phi}=\beta\text{tr}(\rho_{\theta,\phi}H)-S(\rho_{\theta,\phi})\leavevmode\nobreak\ ,$ (10)
where $\text{tr}(\rho_{\theta,\phi}H)$ represents the expected energy of the ansatz, and $S(\rho_{\theta,\phi})$ denotes its Von Neumann entropy. This cost function balances energy minimization and entropy maximization, guiding optimization toward thermal equilibrium. By minimizing $\mathcal{L}_{\theta,\phi}$, the algorithm reconstructs the thermal state, enabling extraction of thermodynamic properties such as energy, entropy, and magnetic susceptibility. More details on obtaining these properties for dinuclear metal complexes using a VQA can be found in Ref.das Neves Silva et al. (2024).
@@ III.2 Ergotropy and Discord Simulation
Figure 2: Simulation test results for different algorithm configurations for four optimizers with 2, 4, and 6 layers considering the average error accumulation define in (a) (left) and the average function evaluations define in (b) (right). As can be seen, L-BFGS-B and COBYLA achieved better results in both metrics.
@@
Figure 3: Simulation of Ergotropy as a function of the temperature using a noiseless and noisy simulator. Simulated noiseless data (filled points) and a noisy simulation (solid line) were obtained from molar magnetic susceptibility calculations, exploring its relationship with ergotropy as described in Eqs. (4). The molar magnetic susceptibility was derived from thermal states approximated by the Variational Quantum Thermalizer (VQT) algorithm, considering temperature values in the range $0\ K-300\ K$ and using 5000 optimization steps with the COBYLA method. The noisy simulation was perform 30 times, with a solid line as its average and the shadow as the standard deviation.
As can be seen, the noiseless simulations show certain consistency with the experimental magnetometry data. The simulation points follow the expected dynamics of ergotropy and discord constants for low temperatures, as expected in theory Cruz et al. (2022). If compared to the experimental data, the algorithm presents a satisfactory accuracy in most of the points, with small oscillations between $\{160\ K-230\ K\}$. This behavior suggests that the algorithm faces intrinsic challenges that lead to discrepancies between simulation results and theoretical or experimental predictions in this particular temperature range.
@@ III.3 Work Extraction Simulation and Protocol Precision
In the following, the protocol for optimal work extraction is applied to the $\rho$ operator, prepared based on ensemble $\{p_{\theta},\left|\psi(\theta^{*},\phi^{*})\right\rangle\}$, considering the regime of pure states for the system for temperatures bellow $83\ K$ Cruz et al. (2022). This state is sent to the protocol, whereby the system is driven from the maximally entangled ground state $\rho_{\left|{\beta_{-}}\right\rangle}$, an active state that stores the maximum amount of energy $E_{0}$, to the passive state $\rho_{\left|00\right\rangle}$, with zero energy stored. This passive state is obtained as an alternative output of the algorithm and is constructed from the states $\left|\psi(\theta^{*},\phi^{*})\right\rangle_{00}$. Considering this system behavior, we apply the protocol for the states with $T<100\ K$ in a noiseless and noisy simulator using the standard configuration, since Fig. 3 shows a maximum ergotropy in this range. The results can be seen as graph bars in Figure 4, which shows the average of probabilities for each state.
The application of the protocol state agrees with the theoretical results for the noiseless simulation, taking all the states to the ground state $\left|00\right\rangle$. For the noisy case, the protocol still shows a probability higher than $80\%$ for the ground state, which shows the low influence of the noise in the process even in room temperatures. Thus, the simulation reveals that the optimal extraction protocol works with accuracy up to $T<100\ K$, although the pure state regime discussed in Cruz et al. (2022) is limited to $T<83K$.
A way to validate the effectiveness of the protocol in the simulation is to see how similar the final state is from the ground state, represented as $\rho_{g}=\left|00\right\rangle\left\langle 00\right|$. Considering this aspect, we compare both using fidelity to quantify the distinguishability (i.e., the degree of similarity) between the state resulting from an optimal work extraction and the output of the protocol. Since $\rho_{\left|00\right\rangle}$ is a pure state, we use $\mathcal{F}(\rho_{\left|00\right\rangle})=\sqrt{\left\langle 0\right|\rho_{\left|00\right\rangle}\left|0\right\rangle}$ for all temperature points. The results can be seen in Figure 5. As expected, $\mathcal{F}(\rho_{\left|00\right\rangle})$ decrease for $T>100\ K$ since the ergotropy decrease in that range.
Figure 5: Fidelity between the simulated state $\rho_{\left|00\right\rangle}$ and the ground state $\rho_{\left|00\right\rangle}$. This metric allows visualizing the protocol efficiency in the extraction of the energy taking a maximally entangled to a ground state configuration. The result shows that the optimal extraction protocol can work with significant accuracy for temperatures up to $100\ K$. The noisy simulation was performed 30 times, with a solid line as its average and the shadow as the standard deviation.
@@
In the noisy case, although the average energy decreases significantly with temperature, the ergotropy does not decrease proportionally. This suggests that noise dissipates energy but maintains a relatively stable fraction of the useful energy extractable as work. Additionally, noise may prevent the system from fully thermalizing, preserving some structure that favors work extraction. Thus, although Ref.Cruz et al. (2022) highlights that the system is capable of energy storage at room temperature, our results reveal a high precision at low temperatures, while the inherent imprecision in the extraction process at higher temperatures could compromise the practical application of these devices. On the other hand, despite the overall degradation of energy efficiency, there is still potential for useful energy conversion even in a noisy environment. This is an important result to characterize the efficiency of the battery even in current quantum hardware standards.
@@ IV Conclusion
Although noiseless simulations closely followed experimental data, simulations in noisy environments exhibited significant deviations as a key factor that degrades battery’s performance. The results show that the presence of environmental noise hinders the accurate evaluation of the energy stored and retrievable from the battery. In this regard, this noise-induced loss of precision highlights the challenge of executing the work extraction protocols on real NISQ hardware without a deeper error mitigation strategy. Moreover, temperature is another critical limitation. While the recent literature underscores the capabilities of the dinuclear battery model based on metal complexes to store energy at room temperature, the work extraction protocol was found to be highly precise only at low temperatures. As the temperature increases, the quantum correlations that enable efficient work extraction are naturally weakened, leading to a noticeable increase in work fluctuation. In fact, in a noiseless scenario, the simulation of the work extraction protocol exhibits high precision only at lower temperatures, with performance degradation observed close to $100\ K$. Conversely, the precision of the work extraction in a noisy environment reveals that fluctuations in quantum correlations can lead to significant deviations, impacting the efficiency of the protocol. These limitations emphasize that, under current conditions, quantum batteries based on dinuclear metal complexes would require careful control of environmental factors or protocol corrections to function reliably.

### v2

@@ III.1 Variational quantum thermalizer and protocol of optimal work extraction
In this work, we employ the Variational Quantum Thermalizer (VQT) algorithm Verdon et al. (2019); das Neves Silva et al. (2024) to approximate the thermal states of the Hamiltonian described in Eq. (1) das Neves Silva et al. (2024). To approximate the thermal states, we utilize a probabilistic quantum ansatz Verdon et al. (2019) $\rho_{\theta,\phi}$, parameterized by variables $\theta$ and $\phi$. These parameters define both the initial state preparation and the structure of the quantum circuit used in the algorithm. Specifically, $\theta$ governs the classical probability distribution $p_{\theta}$ from which computational basis states are sampled, and $\phi$ parameterizes the unitary operations $U(\phi)$ that act on these states via a variational quantum circuit.
In our implementation, schematized in Fig. 1, the block $\hat{V}_{\theta}$ prepares the initial mixed state $\rho^{(i)}_{\theta}$, which encodes a classical probability distribution determined by the parameters $\theta$. This state serves as the input for the variational quantum circuit. The unitary operator $U(\phi)$ is composed of a sequence of parameterized single-qubit rotation gates, including $R_{x}(\phi)$, $R_{y}(\phi)$ and $R_{z}(\phi)$ (rotation around the $x$, $y$ and $z$-axis), along with two-qubit entangling gates such as the controlled-NOT (CNOT). These gates are arranged in alternating layers across the qubits to form the unitary transformation $U(\phi)$, which maps the initial state into a variationally optimized quantum state. The parameter set $\phi$ determines the rotation angles and defines the structure of each variational layer.
The goal of the algorithm is to find optimal parameters $\theta$ and $\phi$ that minimize the quantum relative free energy Verdon et al. (2019) of the ansatz concerning the target thermal state::
$\mathcal{L}_{\theta,\phi}=\beta\text{tr}(\rho_{\theta,\phi}\mathcal{H})-S(\rho_{\theta,\phi})\leavevmode\nobreak\ ,$ (11)
where $\text{tr}(\rho_{\theta,\phi}H)$ represents the expected energy of the ansatz, and $S(\rho_{\theta,\phi})$ denotes its Von Neumann entropy. This cost function balances energy minimization and entropy maximization, guiding optimization toward thermal equilibrium. By minimizing $\mathcal{L}_{\theta,\phi}$, the algorithm reconstructs the thermal state, enabling extraction of thermodynamic properties such as energy, entropy, and magnetic susceptibility. More details on obtaining these properties,as well as implementation details, for dinuclear metal complexes using VQT can be found in Ref.das Neves Silva et al. (2024).
As shown in our previous study, this architecture captures thermodynamic properties of molecular magnetic systems. The clear separation between the latent-state Verdon et al. (2019) preparation governed by $\theta$ and the expressive variational circuit defined by $\phi$ plays a crucial role in both efficiency and physical interpretability of the algorithm.
@@ III.2 Ergotropy and Discord Simulation
Figure 2: Simulation test results for different algorithm configurations for four optimizers with 2, 4, and 6 layers considering the average error accumulation defined in (i) (left-a) and the average function evaluations define in (ii) (right-b). As can be seen, L-BFGS-B and COBYLA achieved better results in both metrics.
@@
Figure 3: Simulation of Ergotropy as a function of the temperature using a noiseless and noisy simulator. Simulated noiseless data (filled points) and a noisy simulation (solid line) were obtained from molar magnetic susceptibility calculations, exploring its relationship with ergotropy as described in Eqs. (5). The molar magnetic susceptibility was derived from thermal states approximated by the Variational Quantum Thermalizer (VQT) algorithm, considering temperature values in the range $0\ K-300\ K$ and using 5000 optimization steps with the COBYLA method. The noisy simulation was perform 30 times, with a solid line as its average and the shadow as the standard deviation. The experimental data was obtained from Cruz et al. (2022), considering the normalized ergotropy
As can be seen, the noiseless simulations show certain consistency with the experimental magnetometry data. The simulation points follow the expected dynamics of ergotropy constant for low temperatures, as expected in theory Cruz et al. (2022). If compared to the experimental data, the algorithm presents a satisfactory accuracy in most of the points, with small oscillations between $\{160\ K-230\ K\}$. This behavior suggests that the algorithm faces intrinsic challenges that lead to discrepancies between simulation results and theoretical or experimental predictions in this particular temperature range.
@@ III.3 Work Extraction Simulation and Protocol Precision
In the following, the protocol for optimal work extraction shown in Fig. 1 is applied to the $\rho$ operator, prepared based on ensemble $\{p_{\theta},\left|\psi(\theta^{*},\phi^{*})\right\rangle\}$, considering the regime of pure states for the system for low temperatures Cruz et al. (2022). This state is sent to the protocol, whereby the system is driven from the maximally entangled ground state $\rho_{\left|{\beta_{-}}\right\rangle}$, an active state that stores the maximum amount of energy $E_{0}$, to the passive state $\rho_{\left|00\right\rangle}$, with zero energy stored. This passive state is obtained as a second output of the algorithm and is constructed from the states $\left|\psi(\theta^{*},\phi^{*})\right\rangle_{00}$. Considering this system behavior, we apply the protocol for the states with $T<100\ K$ in a noiseless and noisy simulator using the standard configuration, since Fig. 3 shows a maximum ergotropy in this range. The results can be seen as graph bars in Figure 4, which shows the average of probabilities for each state. In both cases, the average is taken for all states in the range $T\in[0,100]\ K$ after the protocol application.
The application of the protocol state agrees with the theoretical results for the noiseless simulation, taking all the states to the ground state $\left|00\right\rangle$. For the noisy case, the protocol still shows a probability higher than $80\%$ for the ground state, which shows the low influence of the noise in the quantum information processing, even with the decrease in fidelity at room temperature shown in Fig. 3. Thus, simulation reveals that the optimal extraction protocol works with accuracy up to $T<100\ K$, although the pure state regime discussed in Cruz et al. (2022) is limited to $T<83K$.
A way to validate the effectiveness of the protocol in the simulation is to see how similar the final state is from the ground state, represented as $\rho_{g}=\left|00\right\rangle\left\langle 00\right|$. Considering this aspect, we compare both using fidelity to quantify the distinguishability (i.e., the degree of similarity) between the state resulting from an optimal work extraction and the output of the protocol. Since $\rho_{00}$ is a pure state, we use $\mathcal{F}(\hat{\rho}_{00},\hat{\rho}_{g})=\sqrt{\left\langle 0\right|\hat{\rho}_{00}\left|0\right\rangle}$ for all temperature points. The results can be seen in Figure 5. As expected, $\mathcal{F}(\hat{\rho}_{00},\hat{\rho}_{g})$ decrease for $T>100\ K$ since the ergotropy decrease in that range.
Figure 5: Fidelity between the simulated state $\rho_{00}$ and the ground state $\rho_{g}$. This metric allows visualizing the protocol efficiency in the extraction of the energy taking a maximally entangled to a ground state configuration. The result shows that the optimal extraction protocol can work with significant accuracy for temperatures up to $100\ K$. The noisy simulation was performed 30 times, with a solid line as its average and the shadow as the standard deviation.
@@
In the noisy case, although the average energy has a close correspondence to ergotropy, there is a striking difference in the precision (shadows) — the average energy variance denotes a higher impression. This suggests that noise dissipates energy but maintains a relatively stable fraction of the useful energy extractable as work. Additionally, noise may prevent the system from fully thermalizing, preserving some structure that favors work extraction. Thus, although Ref.Cruz et al. (2022) highlights that the system is capable of energy storage at room temperature, our results reveal a high precision at low temperatures, while the inherent imprecision in the extraction process at higher temperatures could compromise the practical application of these devices. On the other hand, despite the overall degradation of energy efficiency, there is still potential for useful energy conversion even in a noisy environment. This is an important result to characterize the efficiency of the battery even in current quantum hardware standards.
@@ IV Conclusion
Although the experimental results align more closely with the noiseless simulation, the small difference between the noisy and noiseless simulations suggests that the impact of noise in the experiment is minimal for ergotropy, but increase for average energy precision. The results show that the presence of environmental noise hinders the accurate evaluation of the energy stored and retrievable from the battery. Moreover, temperature is another critical limitation. While the recent literature underscores the capabilities of the dinuclear battery model based on metal complexes to store energy at room temperature, the work extraction protocol was found to be highly precise only at low temperatures. As the temperature increases, the quantum correlations that enable efficient work extraction are naturally weakened, leading to a noticeable increase in work fluctuation. In fact, in a noiseless scenario, the simulation of the work extraction protocol exhibits high precision only at lower temperatures, with performance degradation observed close to $100\ K$. Conversely, the precision of the work extraction in a noisy environment reveals that fluctuations in quantum correlations can lead to significant deviations, impacting the efficiency of the protocol. These limitations emphasize that, under current conditions, quantum batteries based on dinuclear metal complexes would require careful control of environmental factors or protocol corrections to function reliably.